

********************************************************************
UARS WEEKLY STATUS REPORT - 15 DEC 2000
********************************************************************
***********
SOLSTICE
***********
This week was a fairly typical one for SOLSTICE in which we collected solar
data on all available orbital tracking periods and stellar calibration
observations whenever possible.
However, on Tuesday (00347) the FOT failed to load our microprocessor loads
per the DAP and the instrument observation list became out of sync with
where the SSPP was pointing. Consequently, at 15:04z on Wednesday (00348)
the instrument was configured for what it thought was a stellar target
when, in fact, the sun was being tracked. This lasted for 8 minutes causing
a sharp degradation in our detectors. The FOT noted some telemetry out of
limits and took the proper, corrective action to safe the instrument
preventing further damage. This has happened in the past but not for a
number of years and the instrument typically makes a slight recovery over
the following few days. The magnitude of this particular incident is under
investigation as we see how the instrument responds over the next week.
Other than this problem, We continue to gather enough data to meet all the
SOLSTICE science objectives during this period of reduced tape recorder
operations.
We are satisfied with the OA_SSPP_ATT patch in regards to how it handles
the data gaps as it has allowed us to process (under RSS) 400 days worth of
data. After analysis of this data we are making some modifications to our
production software.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful operations after powering on December 7, 2000.
All engineering parameters are within nominal operating region and the
science data results are excellent.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now
complete to 9 February 2001.
***********
CDHF
***********
UARS CDHF continues to notify PI Software Representatives regarding Data Gaps
that occur in the LEVEL0 data due to the UARS Spacecraft Anomaly.
DAAC DATA:
- Data created in the month of October, 2000 has been previewed for
HALOE, MLS V5, NMC, PEM, and UKMO in preparation for sending to the DAAC.
UARS CDHF REHOST:
- Need positive response from all PI End Users prior to Rehost transition.
- PI validation of output products and concerns are being addressed.
- Testing of PEM V3.32 (JOBS 1 and 2) are in progress.
- System performance being confirmed.
***********
FOT
***********
REPORTING PERIOD
This report covers 02 December 2000 (Orbit 50431, GMT Day 337) through 08
December 2000 (Orbit 50535, GMT Day 343).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 338/0034 50446 Power On
SUSIM 338/0256 50448 Power On
PEM 339/1317 50469 Day Only
HALOE 341/1320 50499 Power On
A Reverse-to-Forward Yaw Around maneuver (number 94) was successfully executed
in Orbit 50440 beginning 337/14:14:46 and completing at 337/14:51:56 on 02
December 2000. The SOLSTICE instrument was in day only and all other
instruments were in their off or standby mode (including HRDI and PEM). Due to
the SA position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 56 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus voltage was
24.5 volts and the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 22.5 and 23.0 percent respectively. The next (forward-to-
reverse) YAW is scheduled for 31 December 2000 (DOY 366).
The beta angle ranged between the angles of 1.5 degrees to 19.8 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 34 degrees on 17 December 2000 (DOY 352).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 and 1.3 degrees C respectively. Additional battery performance data
for each battery on 1 December 2000 and 8 December 2000 are provided in
Appendix D.
The clock error ranged from Ð10.3 to 12.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 22
¥ PEM = 01
¥ HALOE = 02
¥ ACRIM = 01
¥ SUSIM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 00
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/342 show the observatory data loss to be 300 hours, 01
minutes, 24.888 seconds. This shows an overall decrease of 37.280 seconds since
the last report period. This is a 0.39 percent data loss, which equals a 99.61
percent data capture for the mission.
The decrease of 37.280 seconds occurred over one (1) day:
· 37.280 seconds on Day 339 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS attitude control continues to be the focal point of the FOT's activities
this week. The TAM was placed in update mode on 06 December 2000 for one orbit
(341/1837-2019 GMT). The results of this test showed the TAM performing as
expected with very little increase in the FSS residuals. As a result the TAM
was put into update from 343/1300-2000 GMT on 08 December 2000. This test
yielded much different results. It showed as much as 1.5 degrees of error on
the roll and yaw axes at certain points in the orbit. It was then determined
that this attitude error had to be eliminated for the TAM to work properly on a
converged filter. This will be discussed in detail next week and be
implemented on 12 December 2000.
MEETINGS
The FOT attended a TAM Attitude Meeting on 05 December 2000.
The FOT attended a TAM Attitude Meeting on 06 December 2000.
The FOT attended a monthly UARS Status Meeting on 06 December 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for December 31, 2000 (DOY 366). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 December 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
50439 337/1330 00-686 ACS GYRO BIAS UPDATE FOR FWD FLIGHT
50439 337/1330 00-687 ACS FSS TO MONITOR MODE
50441 337/1603 00-688 ACS FSS TO UPDATE MODE
50446 338/0034 00-689 ACRIM Power ON
50448 338/0256 00-690 SUSIM Power ON - Start
50454 338/1418 00-690 SUSIM Power ON - Complete
50456 338/1606 00-691 OBC Clk Adj 0.00 to -20.6 msec/day
50469 339/1317 00-692 PEM Power ON ZEPS,HEPS day only
50472 339/1834 00-693 OBC Clk Adj -20.6 to 0.00 msec/day
50487 340/1836 00-694 ATC NOOPS D-341
50493 341/0414 00-694 NBTR NOOPS D-341
50499 341/1320 00-695 HALOE Power ON
50501 341/1654 00-697 ACS MAG MODEL TIME COORRECTION
50502 341/1758 00-696 NBTR NOOPS D-341
50502 341/1837 00-698 HALOE Power ON, COMPLETE
50502 341/1837 00-699 ACS TAM TO UPDATE MODE
50503 341/2019 00-699 ACS TAM TO MONITOR MODE
50510 342/0636 00-696 NBTR NOOPS D-342
50517 342/1759 00-700 NBTR NOOPS D-343
50525 343/0639 00-700 NBTR NOOPS D-343
50528 343/1245 00-701 ACS TAM TO UPDATE MODE
50531 343/1652 00-700 NBTR NBTR NOOPS D-343
50532 343/1800 00-702 ACS FWD FLIGHT GYRO BIAS
50532 343/1800 00-703 ACS DUMP TBL 35/36
50532 343/1840 00-704 ACS RESIDUAL LIMITS
50533 343/1940 00-706 ACS FSS
50533 343/1940 00-701 ACS TAM TO MONITOR MODE
50533 343/1940 00-705 NBTR NOOPS D-344
__________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 December 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
__________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 December 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
50456 338/1616 Clk Adj from 0.00 to -20.6 msec/day
50472 339/1843 Clk Adj from -20.6 to 0.00 msec/day
__________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 December 2000
01 Dec 2000 - GMT Day 336 Beta = 03.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------ ------ ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +5.6/-100.8 +5.15/+4.51 +29.6/-13.2 25.1 1.020 21.0
3 +44.8/-11.2 +3.87/+2.92 +28.8/-13.6 25.1 1.020 21.0
08 Dec 2000 - GMT Day 343 Beta = 19.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ --------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 14.2 0.000 0.0
2 +0.0/-112.0 +5.47/+4.51 +30.8/-13.6 25.1 1.040 21.1
3 +22.4/-50.4 +4.51/+3.23 +30.0/-14.8 25.1 1.020 20.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ------ ----------
-no changes V/T 5
_________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 December 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-047 FOT see list Nov 00 14170 Generic Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
SEVNTFW00/49